OpenStudy (anonymous):

A force of 6.00 newtons on a car's bumper makes an angle of 28.5 degrees with the ground. What is the vertical component of this force vector?

OpenStudy (kropot72):

 In the drawing the angle theta is 360 - 28.5 = 331.5 degrees The vertical component of the force vector a is given by \[a _{y}=a \sin \theta =6.00\sin 331.5=you\ can\ calculate\]